13. Chi-square: Which of the following values is significant at the .05 level?

  1. X2(10)=15.43
  2. X2(8)=12.20
  3. X2(5)=11.12
  4. X2(2)=4.23

14. Calculate Test Statistics By Hand

  1. Mean, Median, Mode, Standard Deviation, Variance & Range (use data in the table below)

Athletes’’ weight (lbs)

144

165

210

177

165

182

150

170

190

205

Mean:

Median:

Mode:

Standard deviation:

Variance:

Range:
Answer #1

13.

SNO Caclulated Value P value Significant
a. X2(10)=15.43 0.117153491 NO
b. X2(8)=12.20 0.142500974 NO
c. X2(5)=11.12 0.04905197 Yes
d. X2(2)=4.23 0.12063329 NO

OPTION C is correct.

14.

The sample size is n=10. The provided sample data along with the data required to compute the sample mean and sample variance are shown in the table below:

X X2
144 20736
165 27225
210 44100
177 31329
165 27225
182 33124
150 22500
170 28900
190 36100
205 42025
Sum = 1758 313264

Now, in order to compute the median the data needs to be put into ascending order, as shown in the table below

Position X (Asc. Order)
1 144
2 150
3 165
4 165
5 170
6 177
7 182
8 190
9 205
10 210

Since the sample size n=10 is even, we have that (n+1)/2 = (10+1)/2 = 5.5 is not an integer value, the median is computing by taking the average of the values at the positions 5th and 6th, as shown below

The mode of the data set is 165.

Also, from the information above we get that min=144, max=210, which means that the range is

Range=max−min=210−144=66

Please let me know in comments in case anything is unclear. Will reply ASAP. Do upvote if satisfied!